OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 5x. What is the value of (f o g)(3)? a. -22 b. 3x+8 c. 17 d. 29

OpenStudy (anonymous):

.

OpenStudy (anonymous):

\[(f\circ g)(x) = f(g(x)) \]\[= -2(g(x)) + 8\]\[=-2(5x) + 8\] ... etc.

OpenStudy (anonymous):

And when you're done simplifying, plug in a 3 for the x.